OpenStudy (lin.ivory):

Find the derivative of y=(4x-1)(e^x)(ln(3x+1))

OpenStudy (anonymous):

do you know how to use the chain rule?

OpenStudy (lin.ivory):

i do! ive worked on it for 30 mins with product rule but my answer is different from the key

OpenStudy (anonymous):

could you show your work for me, or do you want me to do it and show my work for you?

OpenStudy (lin.ivory):

can you show your work please? heres what i worked out from before thats apparently incorrect \[y'=(4xe ^{x}+4e ^{x}-e ^{x})(\ln(3x+1))\] \[=(4xe ^{x}+4e ^{x}+4e ^{x}-e ^{x})\ln(3x+1)+(4e ^{x}+4x-1)e ^{x})(\frac{ 3 }{ 3x+1 })\] \[=(4xe ^{x}+7e ^{x})\ln(3x+1)+e ^{x}(4+4x-1)(\frac{ 3 }{ 3x+1 })\] \[=e ^{x}(4x+7)\ln(3x+1)+e ^{x}(4x+3)(\frac{ 3 }{ 3x+1 })\]

OpenStudy (anonymous):

try this: y(x)=f(x)g(x)h(x) product rule with three terms and you get: y'(x)=f'(x)g(x)h(x)+f(x)g'(x)h(x)+f(x)g(x)h'(x)

OpenStudy (anonymous):

sorry for the delay

OpenStudy (lin.ivory):

thats alright!!! THANK YOU SO MUCH!!! :D :D

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

anything else?